OpenStudy (vshiroky):

Write an equation of the line passing through the points (1,10/21) and (3,8/7)

OpenStudy (vshiroky):

@madisonlove

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ \frac{10}{21}}})\quad &({\color{red}{ 3}}\quad ,&{\color{blue}{ \frac{8}{7}}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\) same as before

OpenStudy (anonymous):

y=.9/2x+.2

OpenStudy (vshiroky):

It confuses me with the fractions

OpenStudy (vshiroky):

why is it .9/2+.2